Probability to get exactly 2 heads when flipping a coin 3 times?
The probability of throwing exactly 2 heads in three flips of a coin is 3 in 8, or 0.375.
There are 8 outcomes of flipping a coin 3 times, HHH, HHT, HTH, HTT, THH, THT, TTH, and TTT. Of those outcomes, 3 contain two heads, so the answer is 3 in 8.
First you need to figure out the probability of it not landing on a 2 when you roll one die. Since there are six sides and two sides with 2 on it, the probability that one die will not land on 2 is 4/6.
Now you multiply 4/6 by 4/6 to find the correct odds as to both die not landing on 2.
4/6 * 4/6 = 16/36 = 4/9
The percentage is (4/9)*100% = (approx.) 44.4%
Prob(Sum is odd) = 1/2
Prob(Sum > 6) = 15/36 = 5/12
Prob(Sum is odd or >6) = 27/36 = 3/4
What does the standard deviation tell us about the mean?
The standard deviation tells us nothing about the mean.
Probability of getting all questions wrong on multiple choice test?
In order to answer, the number of questions on the test must be given.
If you roll the die one time, there is a 1/3 probability that you will get a 1 or a 3.
There are 6 possibilities: 1,2,3,4,5,6-----> This number goes on the bottom of your fraction because it is what the probability is out of.
You want 2 different numbers: 1,6--------> This goes on the top because those are the 2 numbers you want.
So the fraction (probability) is 2/6, which simplifies to 1/3.
What is the exponential growth formula?
y = a * bxUsing "^" for powers. Any exponential formula can be written in one of the following two forms:
y = A e^Bx
y = C D^x
where "y" is the independent variable, "x" is the dependent variable, and "A", "B", "C", and "D" are arbitrary constants.
Find the probability that the month of February in a year 2008 may have 5 Tuesdays?
Considering the fact that a non-leap year having 28days in February will definitely have 4 Tuesdays , we have to find the probability that the 29th day of a leap year (2008) is also a Tuesday. This probability is 1/7 i.e 1 out the probable 7days in a week !
Well this said, this question is void in a sense knowing that we are in 2010 and that February 2008 had 5 Fridays and therefore the probability of 5 Tuesdays is ZERO as of 2008 !
How do you find area under a curve between two z values?
Let z1 be the smaller z value and z2 be the larger value. Let N(z) be the cumulative normal distribution evaluated at some value z. The random variable is Z.
The probability that Z has a value from minus infinity to z2 is equal to N(z2). You can show this by drawing a bell shape curve, and shading in everything to the left of z2 as equal to the area under curve.
Similarly, the probability that Z has a value from minus infinity to z1 is N(z1).
The area under the bell curve (standard normal cumulative distribution) is N(z1) - N(z2). I can show this with a little example:
z1= -1 z2 = 2 Area = N(2) - N(-1) = 0.9773 - 0.1587 = 0.8186. I used Excel with the normsdist(z) function. The mean is zero and standard deviation is one with this function.
What does a low mean absolute deviation mean?
It means that the observations are all close to their mean value.
A psychologist wishes to conduct a study on the effects of music deprivation on high school students. A high school class consists of the 30 students numbered in the list below. The researcher establishes a treatment group of 15 students who will have their portable music players replaced by experimental players that present the sound of water running. The control group of 15 students will all get regular portable music players, stuffed full of their favorite songs.
Use the line of random numbers below to select the first six students to receive the treatment. What is the name of the fifth student selected?
59784 44312 15954 09233 00046 74318 02610 57396 16843 38454
Quincy
it is the test one tail
What is the probability that you will spin a 4 and roll a 4?
Assuming each possible number on a spinner has the same probability and an unbiased die is being rolled, the answer depends on how many numbers are on the spinner, and how many times the number 4 appears on each.
To find the probability, workout the probability of spinning a 4 on the spinner and the probability of rolling a 4 on the die; then as spinning the spinner has no effect on rolling the die, they are independent events and to get the probability of both happening multiply them together.
The probability of success is the number of successful outcomes divided by the total number of outcomes, giving:
Probability(spinning a 4) = how_many_4s_are_on_the_spinner / how_many_numbers_are_on_the_spinner
Probability(rolling a 4) = how_many_4s_are_on_the_die / how_many_numbers_are_on_the_die
Probability(spinning a 4 and rolling a 4) = Probability(spinning a 4) × Probability(rolling a 4)
Examples:
Probability values are never negative and are always between 0-1 according to the definition Probability of A= Number of outcomes classified as A/Total number of possible outcomes
Why do population follow normal distribution?
The form of this question incorportates a false premise. The premise is that the data are normally distributed. Actually, is the sample mean which, under certain circumstances, is normally distributed.
What is the difference between Z-Test and The Z-score?
The Z-score is just the score. The Z-test uses the Z-score to compare to the critical value. That is then used to establish if the null hypothesis is refused.
Total number of possible outcomes = 6
Number of successful outcomes = 2
Probability = 2/6 = 1/3 = [ 33 and 1/3rd ] percent.
What is the probability for choosing a red marble?
More information is required. Probability by definition is the proportion of a part, called a sample, to the whole, called a population. Thus in this question, we are given the sample only and without the population, it is impossible to calculate the probability. We need to know the size of the population. As a guide, supposing there are 8 red marbles in a jar containing 40 marbles, then the probability of choosing red is 8/40 or 1: 0.2. There is 20 per cent probability of choosing red.
The probability of a black ball in bag 1 is 4/12 = 1/3.
If you add 3 black balls to bag 2, it will contain 5 black balls out of 15: the probability of a black ball being 3/15 = 1/3.
What is the probability of drawing a red card out of 52?
Since 1/2 of the cards are red, the probability of drawing a red card is 1/2 or 0.5.